OpenStudy (anonymous):

Solve: 1/2 ln (x+3)-ln (x) =0

OpenStudy (anonymous):

ln (x+3)^(1/2)-ln (x) =0 ln (x+3)^(1/2)/x =0 (x+3)^(1/2)/x =1 solve this for x

OpenStudy (anonymous):

How does it equal 1 in the third step?